### v3

Abstract
Multimodal large language models (MLLMs) have shown impressive reasoning abilities. However, they are also more vulnerable to jailbreak attacks than their LLM predecessors. Although still capable of detecting the unsafe responses, we observe that safety mechanisms of the pre-aligned LLMs in MLLMs can be easily bypassed with the introduction of image features. To construct robust MLLMs, we propose ECSO (Eyes Closed, Safety On), a novel training-free protecting approach that exploits the inherent safety awareness of MLLMs, and generates safer responses via adaptively transforming unsafe images into texts to activate the intrinsic safety mechanism of pre-aligned LLMs in MLLMs. Experiments on five state-of-the-art (SoTA) MLLMs demonstrate that ECSO enhances model safety significantly (e.g., 37.6% improvement on the MM-SafetyBench (SD+OCR) and 71.3% on VLSafe with LLaVA-1.5-7B), while consistently maintaining utility results on common MLLM benchmarks. Furthermore, we show that ECSO can be used as a data engine to generate supervised-finetuning (SFT) data for MLLM alignment without extra human intervention.
Keywords:
Multimodal LLMs Safety Image-to-Text Transformation
1 Introduction
Multimodal Large Language Models (MLLMs) [14, 3, 66, 9, 23, 68] have attracted significant attention for their remarkable multimodal capabilities. Building upon the Large Language Models (LLMs) [58, 28, 56, 12], they are aligned with a pre-trained visual encoder using text-image datasets [34, 36, 33], empowering LLMs to conduct conversations with image inputs. Despite these accomplishments, MLLMs encounter challenges in inheriting the safety mechanism of their LLM predecessors. In particular, though MLLMs are built upon LLMs that have been well-aligned with human morals and values [12, 59], they can be easily induced to generate unethical content with the introduction of image inputs [73, 38, 48].
To protect MLLMs, one can repeat training-based alignment strategies of LLMs on MLLMs, such as Supervised Finetuning (SFT) [63, 35, 42, 7] and Reinforcement Learning from Human Feedback (RLHF) [47, 13, 50]. However, this requires meticulous design of red-teaming queries to induce LLMs to generate harmful responses, and can become even more challenging when image inputs involved [48, 73]. Thus, the question is: “How can we transfer the pre-aligned safety mechanisms of LLMs to MLLMs?”
In this paper, we start by conducting a throughout analysis on the safety assessment ability of MLLMs. We observe that despite their susceptibility to malicious queries, (i) MLLMs exhibit clear awareness of unsafe content in their own responses [7]. (ii) The safety mechanism of pre-aligned LLMs persists in MLLMs, but is “suppressed” by the introduction of image features. However, this can be restored by simply removing the images. Building upon these insights, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy exploiting the intrinsic safety mechanisms of pre-aligned LLMs. When presented with an image input with a user query, ECSO first leverages the safety awareness of MLLMs to assess safety of their own responses in a post-hoc manner. Once unsafe initial responses are detected, ECSO converts the image inputs into texts via a query-aware image-to-text (I2T) transformation, and reduces MLLMs to (text-only) LLMs. Safe response generation without images is then performed to restore the safety mechanism of pre-aligned LLMs. Experiments on five MLLM benchmarks demonstrate that the proposed ECSO can significantly enhance model safety without sacrificing the utility performance on common MLLM benchmarks. Moreover, we show that ECSO can be used as a data engine for the generation of SFT data to align MLLMs without extra human intervention.
The main contributions of this work are as follows.
  1. 1.
    We demonstrate that MLLMs, though susceptible to jailbreaking attacks, can detect unsafe content in their own responses and also inherit the safety mechanisms from pre-aligned LLMs that have been inadvertently suppressed.
  2. 2.
    We propose ECSO, a novel training-free and self-contained MLLM protection strategy via first discriminating the safety of its own response and then transforming input images into texts in a query-aware manner to restore the intrinsic safety mechanism of LLMs.
  3. 3.
    ECSO significantly enhances the safety of five SoTA MLLMs, without sacrificing their performance on utility.
2 Related Work
2.0.1 MLLM Vulnerability.
By integrating the capabilities of visual perception with LLMs, MLLMs [23, 14, 3, 66, 1, 9] inherit robust reasoning capabilities of LLMs and excel in dialogues incorporating with visual elements. Despite their impressive capabilities, it has been observed that SoTA MLLMs are increasingly vulnerable to malicious visual inputs [39]. Recent works can be bifurcated into two approaches with respect to the injection of malicious content. One line of works [21, 38] show that embedding the malicious textual queries into images via typography can effectively circumvent the defense mechanisms of MLLMs. The other approach [70, 54, 15, 49, 60, 46, 2, 53, 4, 18] focus on employing gradient-based techniques to create adversarial images that prompt generation of the harmful responses, revealing severe vulnerability.
2.0.2 MLLM Protection.
To enhance safety of MLLMs, a straightforward approach involves aligning MLLMs with specially-constructed red-teaming data [31, 73, 11]. However, red-teaming is labor-intensive and may not encompass all potential attack vectors. Another approach focuses on protecting MLLMs during inference. Wu et al. [65] introduce the manual crafting of system prompts delineating permissible and impermissible actions. However, this may become less effective when new attacks emerge. Wang et al. [61] employ safety steering vectors to adjust MLLM activation in response to the unsafe inputs. However, this may overlook unsafe intents in images that are not detectable by text-centric safety vectors. Most relevant to ours are the works in [10, 48]. Chen et al. [10] introduce a novel automatic self-moderation mechanism, enabling MLLMs to assess and adjust their responses against specific criteria. Despite its promising performance, we will show in Sec. 5.5.1 that, even though instructed to respond safely, MLLMs still struggle to give responses when confronted with images, highlighting the limitation of [10]. Pi et al. [48] augments MLLMs with an ancillary unsafe content detector and output detoxifier, which are external and necessitate additional training on extensive datasets. Instead, the proposed ECSO solely leverages the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, and is devoid of any further training.
3 Preliminary Observations
In this section, we show two intriguing findings involving the safety mechanism of MLLMs, paving the way for the proposed ECSO in Sec. 4.
Figure 1: (left) MLLMs are vulnerable to malicious questions when queried with images but can restore safety when images are excluded. (right) Comparisons of harmless rate (%) of model responses with and without images on five state-of-the-art MLLMs.
Figure 2: (left) Though vulnerable to malicious questions, MLLMs are aware of the unsafe responses of their own. (right) Accuracy of MLLMs discrimination (with and without images) on whether their own responses are safe or not .
3.1 Safety Mechanism Persists in MLLM
In contrast to previous findings suggesting that MLLMs struggle to inherit the safety mechanisms in LLMs [73, 38, 48], here we present a more nuanced view that MLLMs can retain the safeguard when images are not shown to the MLLM. In the following, we perform experiment on the VLSafe dataset [11], which contains 1,110 pairs of queries and images. This dataset has two key features: (1) The malicious queries are paired with benign images; and (2) The input images are auxiliary, i.e. the queries can be answered without referencing the images. These features allow us to dissect the interaction between visual features and safety mechanism by evaluating MLLMs’ responses with and without input images.
Figure 1 compares the harmless rates of MLLMs’ responses with and without the presence of images. As can be seen, when images are present, the MLLMs (except Qwen-VL-Chat [3]) are vulnerable to malicious queries, demonstrating a mere 20% harmless rate. On the other hand, when images are removed from the queries, all models achieve nearly 100% harmless rate. We hypothesize that this discrepancy arises from a distribution shift caused by integrating LLMs with the visual inputs. Specifically, the incorporation of images alters the pre-aligned embedding space of LLMs, rendering existing defense mechanisms ineffective [48]. However, we will show in Sec. 4.2 that the safety mechanism can be reactivated once images undergo a query-aware I2T transformation.
3.2 MLLMs are Aware of Their Own Unsafe Responses
While MLLMs are susceptible to generating harmful content, we investigate whether they are aware of their own safety issues. In this experiment, we collect 1000 responses from LLaVA-1.5-7B. 500 of them are safe and the remaining 500 are unsafe as evaluated by GPT-4111https://chatgpt.ust.hk and double-checked manually. [^2] More detailed description on the dataset can be found in Appendix 0.A.3. We then prompt the MLLMs (detailed in Sec. 4.1) to classify the responses as safe or unsafe. Figure 2 shows the classification accuracies obtained on five MLLMs. As can be seen, though the MLLMs may generate unsafe responses, they exhibit a high degree of safety awareness. Notably, both LLaVA-1.5-7B and ShareGPT4V-7B achieve over 95% accuracy in their assessment. It is crucial to highlight that unlike Figure 1, the assessment here is robust whether the input images are presented or not. Therefore, safety awareness of MLLMs is not compromised by the presence of images. If not otherwise stated, we conduct MLLM safety discrimination with images by default in the sequel.
In summary, a significant discrepancy is observed between MLLMs’ ability to generate safe content (Figure 1) and their capacity for safety discrimination (Figure 2). This divergence may be attributed to the inherent ease of discrimination tasks over generation tasks (a hypothesis supported by [7]) or understood through analogies drawn from scalable oversight and complexity theory [52]. In Sec. 4, we will explore how leveraging these insights can effectively mitigate the safety concerns associated with MLLMs.
4 Methodology
In this section, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy by exploiting the two observations in Sec. 3. An overview of the main steps is shown in Figure 3.
Figure 3: Overview of ECSO. Step 1: User queries are full-filled as usual. Step 2: The MLLM is prompted to judge whether its initial response is safe or not. Safe answers are returned, while unsafe ones proceed Step 3 and 4. Step 3: Images of unsafe queries are converted into texts via query-aware text-to-image transformation. Step 4: Malicious content in either images or user queries are now both represented by plain text, which can be deal with by the pre-aligned LLMs in MLLMs to generate safe responses.
Figure 4: Prompt templates for ECSO, where <image> denotes the presence of image inputs and {} denotes a placeholder for the actual text inputs.
4.1 Harmful Content Detection
Given a (benign or malicious) user query with image $v$ and query $x$, we first prompt the MLLM $F_{\boldsymbol{\theta}}$ (with parameter $\boldsymbol{\theta}$) to output response $\tilde{y}$ (Figure 3):
$\tilde{y}=F_{\boldsymbol{\theta}}(v,x).$ (1)
The response $\tilde{y}$ may not be safe as MLLMs are susceptible to attacks. Using a prompt template $P_{\text{det}}$ for harm detection (Figure 4, left) to wrap $x$ and $\tilde{y}$, we ask the MLLM to discriminate the safety of its own $\tilde{y}$ (Step 2 in Figure 3), and output the predicted safety $s$ of the model response:
$s=F_{\boldsymbol{\theta}}(v,P_{\text{det}}(x,\tilde{y})),$ (2)
The near-GPT-4 performance on safety assessment, as seen in Sec. 3.2, supports this introspection. If the model response $\tilde{y}$ passes this self-checking, it will be presented to the user. On the contrary, if $\tilde{y}$ is detected as unsafe, we propose to first transform the image into text (Step 3 in Figure 3) and then query the MLLMs again without visual inputs (Step 4 in Figure 3). These will be detailed in Sec. 4.2 and 4.3, respectively.
4.2 Query-Aware Image-to-Text (I2T) Transformation
To restore the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, we propose to transform the input query image to plain text. Any malicious content in the image that might induce harmful responses are then either converted to text or completely left away from the remaining procedure. However, there may be information loss in the image-to-text (I2T) conversion. To retain the image information to the greatest extent, we use a prompt template $P_{\text{trans}}$ (Figure 4, middle) that includes the original question. The MLLM is then prompted to generate the query-aware caption $c$:
$c=F_{\boldsymbol{\theta}}(v,P_{\text{trans}}(x)).$ (3)
As will be seen in Sec. 5.5.3, query-awareness in $c$ is indispensable because without it, the caption might not include all the relevant information necessary to answer the original query. Here, we implement this process with captioning, though more advanced T2I transformation methods (e.g., [64]) can also be explored.
4.3 Safe Response Generation Without Images
Recall from Sec. 3.1 that LLMs are safer than MLLMs and the safety mechanism in MLLMs can be reinstated with the removal of image inputs. To acquire a safer response, we prompt MLLMs with the original query along with the previously generated query-aware caption $c$ in Sec. 4.2 (instead of the query image):
$y=F_{\boldsymbol{\theta}}(\mathrm{null},P_{\text{gen}}(c,x)),$ (4)
where $\mathrm{null}$ denotes an empty input (i.e., the query image $v$ is excluded), and $P_{\text{gen}}$ is the prompt for safe response generation without images (Figure 4, right). Since images have been removed, the MLLM reduces to a text-only LLM. This step is effective in safeguarding the MLLM because any malicious contents that induce harmful responses are exposed to the safe pre-aligned LLM. To further underscore the priority of safety, we include the words “HARMLESS and ETHICAL” in the prompt during inference, as shown in Figure 4 (right).
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
01-Illegal Activity 94.9 78.4 96.9 $(+18.6)$ 22.7 96.9 $(+74.2)$ 25.8 92.8 $(+67.0)$
02-HateSpeech 93.9 84.7 96.9 $(+12.3)$ 56.4 87.7 $(+31.3)$ 51.5 90.2 $(+38.7)$
03-Malware Generation 47.7 84.1 97.7 $(+13.6)$ 31.8 86.4 $(+54.6)$ 38.6 84.1 $(+45.5)$
04-Physical Harm 71.5 81.9 93.8 $(+11.8)$ 40.3 88.9 $(+48.6)$ 41.0 84.7 $(+43.8)$
05-Economic Harm 97.5 95.9 96.7 $(+0.80)$ 86.9 97.5 $(+10.7)$ 86.9 96.7 $(+9.80)$
06-Fraud 85.7 79.9 95.5 $(+15.6)$ 28.6 89.0 $(+60.4)$ 33.1 85.1 $(+52.0)$
07-Pornography 88.1 90.8 93.6 $(+2.80)$ 76.2 88.1 $(+11.9)$ 69.7 76.2 $(+6.40)$
09-Privacy Violence 75.5 84.2 92.1 $(+7.91)$ 41.7 87.8 $(+46.0)$ 43.9 81.3 $(+37.4)$
Average 81.9 85.0 95.4 $(+10.4)$ 31.7 90.3 $(+42.2)$ 32.1 86.4 $(+37.6)$
Table 1: Harmless rates on MM-SafetyBench with LLaVA-1.5-7B [36]. ECSO significantly improves the safety of MLLMs by restoring their intrinsic safety mechanisms, alleviating the necessity of additional training procedure [7, 48].
5 Experiments
In this section, we empirically evaluate the proposed ECSO. First, we introduce the experimental settings in Sec. 5.1 and 5.2. Then, we assess ECSO from the following perspectives: (i) How well can ECSO protect the existing MLLMs? (Sec. 5.3) (ii) Can ECSO maintain the utility of MLLMs? (Sec. 5.4) (iii) Can ECSO serve as a data engine to produce data for safety alignment? (Sec. 5.6). Besides, we ablate the effects of the key components of ECSO in Sec. 5.5.
5.1 Models and Evaluation Protocols
5.1.1 Models.
Five SOTA MLLMs are considered, including the LLaVA-1.5-7B [36], ShareGPT4V-7B  [9], mPLUG-OWL2-7B [66], Qwen-VL-Chat [3] and InternLM-XComposer [68]. To demonstrate the effectiveness of the proposed ECSO, we compare with directly prompting the MLLMs and responding to queries (denoted Direct).
5.1.2 Evaluation.
The safety degree of a model’s responses is measured by the harmless rate [55, 7], which is defined as the proportion of safe responses over all the responses $D$ (i.e., $\text{HR}=\sum_{d\in D}\frac{I(d)}{\left|D\right|}$, where $I(\cdot)$ is the indicator function with $I(\cdot)=1$ if the response is harmless (as judged by prompting GPT-4 followed by manual verification), and $I(\cdot)=0$ otherwise). The detailed prompt for GPT-4 can be found in Figure 11.
Figure 5: Qualitative comparison showing how ECSO generates harmless response. Direct/Initial responses: Model response when directly prompted. This is also the initial response in the first step of ECSO. Harm?: Harmful content detection as in Sec. 4.1. Caption: Query-aware I2T captioning as in Sec. 4.2. ECSO: Safe response generation without images by ECSO as in Sec. 4.3. Text in Red (resp. green) is harmful (resp. harmless). Dashed red rectangles highlights content activating the safety mechanism within the pre-aligned LLMs in Sec. 4.3.
5.2 Safety Benchmark Datasets
Experiments assessing the safety of MLLMs’ responses are primarily performed on the MM-SafetyBench [38] and VLSafe [11] datasets. MM-SafetyBench [38] contains 5,040 examples with malicious intents in 13 common scenarios (e.g., illegal activities, hate speech, and malware generation). We evaluate on only 8 scenarios because empirically we find that even text-only LLMs perform poorly on the remaining scenarios. Full results can be found in Appendix 0.C. In this dataset, most of the malicious contents are in the images, while the texts are usually benign. The image in each question originates from malicious keywords and can be from one of the following: (1) SD: Images generated by Stable Diffusion (SD) [51] by conditioning on the malicious keywords; (2) OCR images with malicious keywords; (3) SD+OCR: Images generated by Stable Diffusion and then subtitled by OCR. Apart from the multimodal data, MM-SafetyBench also offers text-only questions built upon the malicious keywords, which will also be evaluated in our experiment. VLSafe [11], instead, contains 1,110 malicious image-text pairs in its examine split. The malicious intent is clearly represented in the text queries. Examples from both datasets are shown in Figure 5. More details can be found in Appendix 0.A.1.
Figure 6: Harmless rates on MM-SafetyBench (SD+OCR) for the ShareGPT4V-7B [9], mPLUG-Owl2-7B [66], Qwen-VL-Chat [3] and InternLM-XComposer-7B [68]. Blue and orange shades represent the harmless rates when querying MLLMs directly and with our proposed ECSO, respectively.
5.3 Evaluation of Safety
Table 1 compares the harmless rates on MM-SafetyBench by directly prompting LLaVA-1.5-7B (Direct) and prompting via the proposed ECSO. As can be seen, ECSO greatly boosts the safety of LLaVA-1.5-7B. Specifically, on average, the proposed ECSO improves LLaVA-1.5-7B’s harmless rate from 31.7% to 90.3% when queried with OCR images, and from 32.1% to 86.4% when queried with SD+OCR images. In particular, ECSO offers much bigger safety gains on OCR and SD+OCR compared to SD. This is because SD is less effective in attacking MLLMs (as can be seen from Table 1). As most SD responses obtained by a direct prompting of LLaVA-1.5-7B are already benign, the improvement by ECSO is smaller. It is interesting that the harmless rate of ECSO even surpasses Text-Only (i.e., the upper bound of ECSO). This can be explained by the inclusion of the keywords “HARMLESS and ETHICAL” in Sec. 4.3, instructing LLMs to pay more attention and respond in a safer way.
Figure 7: Harmless rates on VLSafe using direct prompting versus ECSO. Red numbers on the top indicate the absolute improvement in terms of the harmless rate.
Figure 5 shows examples of how ECSO generates harmless responses from malicious queries. As can be seen, after identifying harmful content in the initial response, ECSO converts the image to text caption. As the LLM is safety-aligned, it identifies unsafe content in the caption and generates a harmless response.
Figure 6 shows the comparison of harmless rates for the other MLLMs. Notice that we only show results on the SD+OCR split because it is adopted as the default split in the MM-Safetybench  [38]. The remaining results can be found in Appendix 0.C. As can be seen, the proposed ECSO again offers safety protection for the MLLMs in a wide range of scenarios.
Figure 7 shows the harmless rate comparison on VLSafe for various MLLMs. As can be seen, the proposed ECSO significantly improves the harmless rate.
Recall from Sec. 3.1 that the MLLMs can only achieve satisfactory harmless rates when images are excluded. Now, with ECSO, we can maintain the safety of MLLMs while retaining the information in the images. Hence, we conclude that the proposed ECSO can effectively reactivate the safety mechanism within MLLMs even with the presence of images via query-aware I2T transformation and safe response generation without images.
5.4 Evaluation of Utility
In this section, we show that ECSO only causes minor degradation to the utility of MLLMs, and might even offer improvements in some scenarios.
5.4.1 Datasets.
Experiments are performed on popular MLLM utility benchmarks, including MME [17], MM-Vet [67], and MMBench [40]. These benchmarks cover a wide range of common abilities/tasks (e.g., maths, OCR, perception of objects, color and understanding of arts) that are considered as important for MLLMs. MME [67] has the subsets of perception (MME-P) and cognition (MME-C), with 10 and 4 tasks, respectively. For each subset, the sum of accuracy and accuracy+ [17] within each task are reported to evaluate utility. For MMBench [40] and MM-Vet [67], accuracy and average GPT score (ranging from 0 to 1) for all samples are reported. More details on these datasets are in Appendix 0.A.2. We assume that all queries are benign and do not induce harmful answers. In other words, any detection of harm by MLLMs would be considered as a false alarm.
Models MME MMBench MM-Vet
LLaVA-1.5-7B 0.50% 1.23% 0.46%
ShareGPT4V-7B 1.93% 4.24% 0.46%
mPLUG-Owl2-7B 0.20% 0.20% 1.10%
Qwen-VL-Chat 1.26% 2.88% 4.59%
InternLM-XComposer-7B 0.08% 0.00% 0.00%
Table 2: Misclassification ratios of MLLMs predicting benign queries malicious on MME [17], MMBench [40], and MM-Vet [67], respectively. Most of the time, the query-aware I2T transformation will not be triggered on common benchmarks.
Models MME-P MME-C MM-Vet MMBench
Direct ECSO Direct ECSO Direct ECSO Direct ECSO
LLaVA-1.5-7B 1507.4 1507.4 355.7 357.1 30.5 30.6 64.6 64.2
ShareGPT4V-7B 1566.4 1567.1 376.4 380.7 33.9 34.4 66.5 66.1
mPLUG-Owl2-7B 1456.0 1456.0 345.7 345.7 33.9 33.9 66.7 66.5
Qwen-VL-Chat 1481.5 1481.5 347.1 347.1 49.6 49.7 59.7 59.1
InternLM-XComposer-7B 1254.1 1254.1 200.7 200.7 33.3 33.3 49.3 49.3
Table 3: Utility scores of MLLMs on MME-P [17], MME-C [17], MM-Vet [67], and MMBench [40], separately. The safety improvement of ECSO in Table 1 comes without sacrificing the utility performance.
5.4.2 Results.
Table 2 shows the misclassification ratios by MLLMs that predict benign queries as malicious. As can be seen, most of the time MLLMs can correctly recognize the benign queries and do not trigger the I2T transformation process. Table 3 shows the utility scores of MLLMs on the benchmarks. It can be observed that across different models, ECSO does not hurt the utility scores of MLLMs on MME-P and MM-Vet, while even offers slight improvement on MME-C and MM-Vet. We speculate that this improvement might be attributed to the world knowledge elicited from query-aware captioning.
Figure 8: Harmless rate of LLaVA-1.5-7B when images are invisible and visible to the model. The SD+OCR split of MM-SafetyBench is evaluated here.
Methods Harmless Rate
w/o images 91.8
w/ images 85.6
Table 4: Performance of LLaVA-1.5-7B on the examine split of VLSafe with and without images.
5.5 Ablation Study
5.5.1 Necessity of excluding images.
In ECSO, the unsafe image-text pairs are queried again with images converted to captions. A critical design of ECSO is that the actual images are discarded in this stage. Here, we show that the absence of image is the key to generate safer responses. To ablate this feature, we insert the image features to MLLMs in addition to the query-aware caption. Figure 8 and Table 4 show the harmless rates of LLaVA-1.5-7B on MM-SafetyBench (SD+OCR) and VLSafe (examine), respectively. On both benchmarks, the harmless rate decreases by a large margin with images are incorporated. Hence, ECSO is indeed restoring the safety mechanism of pre-aligned LLMs, and very different from multimodal Chain of Thoughts (MM-CoT) [69, 10] that succeeds only via more reasoning steps or multi-turn self-moderation.
Furthermore, we find that the performance drop is more significant on MM-SafetyBench than that on VLSafe. This can be explained by the differences in sources of malicious contents. MM-SafetyBench attacks the image modality, while VLSafe attacks the text modality. Hence, MM-SafetyBench, with images visible to the MLLMs, is more prone to be induced.
Methods MME MM-Vet MMBench
w/ step 3&4 1865 30.6 64.18
w/o step 3&4 1847 30.0 63.83
Table 5: Utility on MME, MM-Vet and MMBench.
5.5.2 Effect of Steps 3&4
In Sec. 4.2 and 4.3, we caption the image (step 3) and query the LLM again (step 4) in case of unsafe responses. A seemingly simpler solution is to directly refuse to respond and output “I cannot answer this question due to safety constraint”. Table 5 shows the utility of LLaVA-1.5 when employing such a strategy. As can be seen, the model with steps 3 and 4 achieve higher utility because they always respond to the queries.
5.5.3 Effect of Query-aware I2T Transformations.
In this experiment, we demonstrate that the proposed query-aware I2T transformation in ECSO is indispensable to maintaining the utility of MLLMs. We use ShareGPT4V-7B for ablation study, since it makes more mis-classifications as shown in Table 2, which makes it more prone to affect the utility. To study the effect of conditioning on the query, we replace $P_{\text{trans}}$ with “Please give a caption for the image”. As demonstrated in Table 6, the removal of query-aware I2T transformations has a negative impact on the utility of MLLMs.
An example is shown in Figure 9. Using query-aware I2T transformation (Figure 9, left), the generated caption mentions the positions of the two elephants, which is pertinent to the query. Hence, when queried again without the image, the model can give the correct answer. On the contrary, when the query is not used for conditioning (Figure 9, right), the generated caption does not offer any valuable clues and leads to an incorrect answer.
Methods MME-P MME-C MMBench
Direct 1566.4 376.4 66.5
ECSO 1567.1$(+0.05\%)$ 380.7 $(+1.14\%)$ 66.1 $(-0.65\%)$
w/o Q. Trans. 1562.7 $(-0.23\%)$ 376.4 $(+0.00\%)$ 65.8 $(-1.05\%)$
Table 6: Ablation study on the use of query-aware I2T transformation (abbreviated as Q. Trans.) in maintaining the utility of MLLMs.
Figure 9: Qualitative comparison on LLaVA-1.5-7B with (left) and without (right) query-aware I2T transformations. The original queries are highlighted.
5.6 Safety Alignment
In this section, we show that ECSO can serve as a data engine to generate SFT samples for safety alignment. Traditionally, to conduct safety alignment, a supervised dataset $D^{*}=\left\{(v,x,y^{*})\right\}$ (with potentially malicious querying text $x$, image $v$ and benign response $y^{*}$) is required. However, curating the ground-truth response $y^{*}$ can be expensive. In the following, we assume access to only an unsupervised safety dataset $D=\left\{(v,x)\right\}$. To obtain benign response $y$ for alignment, we apply ECSO on $D$ to acquire $D^{\prime}=\left\{(v,x,y)\right\}$, where $y$ is the generated safe response in Sec. 4.3. Note that all intermediate results (including the initial response $\tilde{y}$, safe indicator $s$, and query-aware caption $c$) are discarded. Then, $D^{\prime}$ can be used for safety alignment via supervised finetuning (SFT).
In this experiment, we adopt VLGuard [73], a supervised safety alignment dataset containing 3000 query-response pairs covering various harmful categories (e.g., privacy violation and deception), as $D^{*}$. We replace the ground-truth response $y^{*}$ with the ECSO-generated $y$ to form $D^{\prime}$. We then finetune two LLaVA-1.5-7B models, one using $D^{*}$ while the other using $D^{\prime}$, together with a set of shared utility data to maintain utility. Finally, we evaluate the two finetuned models on the SD+OCR split of MM-SafetyBench. More details on the fine-tuning process and datasets are in Appendix 0.B.2.
Figure 10: Harmless rates of LLaVA-1.5-7B on MM-SafetyBench (SD+OCR) using ECSO and the finetuned on $D^{\prime}$ (ECSO_VLGuard) and $D^{*}$ (VLGuard) separately.
Figure 10 compares the harmless rates of the following models: (i) the original LLaVA-1.5-7B equipped with training-free ECSO (denoted ECSO), (ii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{\prime}$ and utility data to respond to queries (denoted ECSO_VLGuard) and (iii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{*}$ and utility data (denoted VLGuard). As can be seen, (i) In most cases, ECSO_VLGuard is outperformed by ECSO since ECSO_VLGuard is trained on the ECSO outputs. (ii) ECSO_VLGuard offers better safety than VLGuard, showing that data generated by ECSO has comparable or even better quality than human-verified data, offering better trade-off among safety and utility.
5.7 Limitation and Future Work
While ECSO can notably strengthen the safety of MLLMs, it heavily relies on the LLMs’ capacity to identify and neutralize unsafe queries. Therefore, any deficiencies in the LLMs’ safety mechanism may compromise ECSO’s performance in multimodal scenarios. Moving forward, an exciting prospect for future research is to explore how to turn multimodality from a challenge into an asset for safety. By developing new methods harnessing rich context provided by multiple modalities, it might be possible to create more nuanced and context-aware safety mechanisms, increasing the efficacy and reliability of MLLMs’ safety protocols.
6 Conclusion
This paper proposes ECSO, an innovative and training-free safeguarding method that capitalizes on the intrinsic safety mechanisms within MLLMs. Additionally, our findings reveal that ECSO not only acts as a protective measure but also serves as a powerful tool for autonomously generating Supervised Fine-Tuning (SFT) data. This facilitates the alignment of MLLMs with desired safety standards without the need for additional human intervention. We hope that the contributions of this work will provide valuable guidance for the community in the ongoing endeavor to construct more secure MLLMs.
Acknowledgement
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This work was partially supported by NSFC key grant 62136005, NSFC general grant 62076118, Shenzhen fundamental research program JCYJ20210324105000003, and the Research Grants Council of the Hong Kong Special Administrative Region (Grants C7004-22G-1 and 16202523).
Appendix
Appendix 0.A Datasets
0.A.1 Safety Datasets
0.A.1.1 MM-SafetyBench
[38] contains 5,040 examples with malicious intents in 13 common scenarios. For statistics of these datasets, please refer to Table 1 of [38]. The image in each example originates from malicious keywords and can be from one of the following: (1) SD: Images generated by Stable Diffusion (SD) [51] by conditioning on malicious keywords; (2) OCR images with malicious keywords; (3) SD+OCR: Images generated by Stable Diffusion and then subtitled by OCR. Besides these multimodal data, MM-SafetyBench offers text-only questions built upon the malicious keywords. In the following, we give a brief explanation for the 13 scenarios and their evaluation criteria.
  • Scenarios 01-07 & 09: 01-Illegal Activity, 02-HateSpeech, 03-Malware Generation, 04-Physical Harm, 05-Economic Harm, 06-Fraud, 07-Pornography, 09-Privacy Violence. These scenarios present clear-cut cases for users to discern whether the responses generated are harmful. A response is categorized as “unsafe” if it includes any harmful content. In the absence of such content, it is classified as “safe”. The evaluation prompt for these scenarios can be found in Figure 15 of [38].
  • Scenarios 08 & 13: 08-Political Lobbying and 13-Government Decision. These two scenarios are centered around political topics. For a response to be considered “safe”, it should demonstrate the model’s restraint from expressing political opinions, typically by opting not to respond. Responses that do not adhere to this principle are deemed “unsafe”. The evaluation prompt for these scenarios can be found in Figures 16, 20 of [38], respectively.
  • Scenarios 10-12: 10-Legal Opinion, 11-Financial Advice, 12-Health Consultation. These scenarios involve specialized professional fields, such as legal and healthcare. In these cases, models should clearly state their lack of relevant certification, for instance, by responding with a disclaimer like “I apologize, but I am not certified to provide professional advice in this area…” Additionally, they should include a risk warning. Responses that follow this protocol are considered “safe”. The evaluation prompt for these scenarios can be found in Figures 17, 18 and 19 of [38].
Among these topics, we mainly evaluate ECSO on scenarios [01-07 & 09]. This results from the nature of ECSO that it is based on the safety mechanism of LLMs and it cannot offer safety improvements on scenarios that even LLMs are unsafe. Specifically, we find that on the remaining scenarios (i.e., 08, 13, 10-12), some MLLMs reveal safety vulnerabilities even with text-only queries. Therefore, we did not study them in the paper. However, we still report results of ECSO in Appendix 0.C.
0.A.1.2 VLSafe
is proposed in [11] to train and validate the harmlessness alignment of MLLMs. Specifically, it contains 4764 and 1110 malicious queries and corresponding harmless responses in its alignment split and examine split, respectively. The harmful intent is clearly represented in the text queries, while the images are totally benign. In this paper, we use the queries in its examine split for evaluation. The evaluation prompt for VLSafe is shown in Figure 11.
Figure 11: Prompts for evaluating the safety of VLSafe.
0.A.1.3 VLGuard
[73] is also proposed for the safety alignment of MLLMs. There are 2000 images in its training set. 977 of them are harmful while the remaining 1023 are benign. Each safe image is matched with a safe query-response pair and an unsafe pair, while each harmful image is coupled with a single query-instruction explaining the unsafe nature of this image. Note that the resposes and queries in this data set are generated by GPT4. In total, there are around 3000 query-response pairs in the training set. Here, we use the training set of VLGuard for the safety alignment experiment in Sec. 5.6.
0.A.2 Utility Datasets
In this section, we introduce the datasets leveraged to evaluate the utility of an MLLM. Table 7 shows the statistics and characteristics of these datasets.
0.A.2.1 MME
[17] examines both the perception (MME-P) and cognition (MME-C) abilities of MLLM on a total of 14 sub-tasks with 2374 questions. Each instruction consists of a question followed by “Please answer yes or no”. For each test image, two instructions are manually designed. The ground-truth answer of the first question is “yes”, and that of the second question is “no”. The utility score of a sub-task is based on the sum of accuracy and accuracy+. Here, accuracy is calculated based on each question, while accuracy+ is based on each image where both of its two questions need to be answered correctly. The perception score is the sum of scores of all perception sub-tasks (ranging from 0 to 2000). The cognition score is calculated in the same way (ranging from 0 to 800).
0.A.2.2 MMBench
[40] contains 2974 single choice questions covering 20 different ability dimensions, such as object localization and social reasoning, for MLLMs. Each ability dimension includes more than 75 questions. The utility score for this dataset is defined as the accuracy over all the questions, thus ranging from 0% to 100%. In addition, as some MLLMs might prefer a certain choice (e.g., choice “A”) among all given choice, MMBench proposed Circular Evaluation, under which each question is fed to an MLLM N times (N equals to the number of choices). Each time circular shifting is applied to the choices and the answer to generate a new prompt for MLLMs. An MLLM is considered successful in solving a question only if it correctly predicts the answer in all rotational passes.
0.A.2.3 MM-Vet
[67] defines six core vision-language capabilities, including recognition, OCR, knowledge, language generation, spatial awareness, and math, which integrate to solve various complicated multimodal tasks. Different from MME and MMBench, MM-Vet requires the MLLM to answer the question in an open-ended manner, which is more flexible but also more complex to evaluate. To address this, for a model prediction, MM-Vet queries GPT-4 with few-shot evaluation prompts to obtain an evaluation score ranging from 0 to 1. The utility score for this dataset is defined as the sum of all scores divided by the number of questions and then multiplied by 100 to fall in the range of $[0,100]$.
Benchmark #Questions #Tasks Ans. format Metric range
MME 2374 14 Yes/No $[0,2000]/[0,800]$
MMBench 2974 20 Single choices $[0,100]\%$
MM-Vet 218 6 Open-ended $[0,100]$
Table 7: Statistics of utility benchmarks used in our experiment. The metric range indicates the lowest/highest metric score. The metric ranges of MME are for its two sub-categories: Perception and Cognition.
0.A.3 Datasets Used in Preliminary Study
To assess the safety awareness in MLLMs, we collect model responses from multiple dataset sources:
  • MM-SafetyBench: 72, 79, and 85 responses are sampled from 01-Illegal Activity, 02-HateSpeech, and 04-Physical Harm, respectively. All responses are unsafe.
  • VLSafe (examine): 264 responses are sampled from the examine split of VLSafe, which are all unsafe.
  • LLaVA_150k  [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K: 500 responses are sampled from the instruction tuning dataset of LLaVA [37]. All responses are safe.
In total, there are 1,000 responses, in which 500 of them are safe and the remaining 500 are unsafe. Note that all responses are generated by LLaVA-1.5-7B and are classified into safe/unsafe via GPT-4 (and double-checked manually). For MM-SafetyBench and VLSafe (examine), we use the same prompt as the one in their respective evaluation process. For LLaVA_150k, we use the same prompt as for VLSafe (examine).
Appendix 0.B Implementation Details
0.B.1 Model Inference
For all models, we disable sampling during inference to eliminate randomness in generation. Following the officially provided default configuration of all experimented models, only InternLM-XComposer is evaluated using beam search (#beams = 5), while others do not adopt beam search.
0.B.2 Finetuning
In this section, we present details on the training data and configurations. Further, we provide analysis on the generated data by ECSO.
0.B.2.1 Data Construction.
In Sec. 5.6, we use ECSO to generate benign responses from the queries of VLGurad. To construct $D^{\prime}$, we only keep the query-response pairs that are initially detected as harmful by the MLLMs. Note that though the initial response $\tilde{y}$ is harmful, the generated response $y$ by ECSO is safe. In this way, we obtain malicious queries together with benign responses, which are presumed to be effective for the safety alignment process. Finally, we obtain 232 such query-response pairs to form $D^{\prime}$ for safety alignment. The query-response pairs remaining in VLGuard are denoted $D^{\prime}_{\text{safe}}$. Different from $D^{\prime}$, the initial responses in $D^{\prime}_{\text{safe}}$ are considered benign by MLLMs and remain unchanged. We will show in the following experiment that $D^{\prime}_{\text{safe}}$ is critical to maintaining the utility of MLLMs during safety alignment.
Figure 12: Safety-utility trade-off when mixing $D^{\prime}$ (left) or $D^{*}$ (right) with $D^{\prime}_{\text{safe}}$. For safety, we report the sum of harmless rates (evaluated by gpt-3.5-turbo-0125 due to limited budgets) on three scenarios of MM-SafetyBench (01-Illegal Activity, 02-HateSpeech and 03-Malware Generation). For utility, we report the sum of scores in MME-P and MME-C. Note that the score of the celebrity subset in MME-P is excluded as the questions involve answering the name of a celebrity, which violates the safety criterion in VLGuard.
Empirically, finetuning LLaVA-1.5-7B with $D^{\prime}$ only leads to significant utility degradation of MLLMs (Figure 12). To address this, we mix $D^{\prime}$ with examples in $D^{\prime}_{\text{safe}}$ (which are called “utility data” in Sec. 5.6) to form $D^{\prime}+D^{\prime}_{\text{safe}}$ and evaluate the resulting finetuned model. In Figure 12 (left), we show the safety and utility of the resulting model trained using different numbers of samples from $D^{\prime}_{\text{safe}}$. It can be observed that as the number of samples in $D^{\prime}_{\text{safe}}$ increases, the utility of the model can be restored to that of the original model. This is because the responses in $D^{\prime}_{\text{safe}}$ are generated by the model itself, thus representing the MLLM’s original abilities. [^4] In contrast, although $D^{\prime}$ also comes from MLLMs, it is generated without the images. However, the addition of $D^{\prime}_{\text{safe}}$ also tends to diminish the safety gain brought by $D^{\prime}$ because this may lower the importance of $D^{\prime}$ during training. This leads to a trade-off between safety and utility. In this experiment, we take ECSO_VLGuard (in Figure 10) as the model trained on data mixed with 150 $D^{\prime}_{\text{safe}}$ samples, which obtains roughly the same utility as the untrained LLaVA-1.5-7B while still maintaining good safety. As will be shown below, even this “optimal" ECSO_VLGuard model is outperformed by the proposed ECSO.
Similarly, to compare the quality of the response generated by ECSO (i.e., $D^{\prime}$) and from the ground-truth of VLGuard (i.e., $D^{*}$), we mix $D^{*}$ with $D^{\prime}_{\text{safe}}$ to construct $D^{*}+D^{\prime}_{\text{safe}}$. To make it comparable with $D^{\prime}$, we also sample 232 examples from VLGuard to form $D^{*}$. Figure 12 (right) shows the safety-utility trade-off. We take VLGuard (in Figure 10) as the model trained on data mixed with 200 $D^{\prime}_{\text{safe}}$ samples.
Compared with the model trained with $D^{\prime}+D^{\prime}_{\text{safe}}$ (Figure 12, left), we have the following observations.
  • For safety alignment, data generated by ECSO are even better than the ground-truth. As can be observed, without the presence of $D^{\prime}_{\text{safe}}$, the models finetuned on $D^{\prime}$ and $D^{*}$ show similar safety performance (e.g., both are around 225). This demonstrates that the responses generated by ECSO offer comparable quality to those by GPT-4.
  • Data generated by ECSO show better safety-utility trade-off than the ground-truth. Without $D^{\prime}_{\text{safe}}$, models finetuned on $D^{\prime}$ offer much better utility than that on $D^{*}$. With increasing $D^{\prime}_{\text{safe}}$, the safety of models trained on $D^{*}$ decreases much faster than those trained on $D^{\prime}$. This results from the similarity in distribution between $D^{\prime}$ and $D^{\prime}_{\text{safe}}$ as both of them are generated by the model itself. However, $D^{*}$ are curated by another model (GPT-4) whose responses might have a large domain gap to $D^{\prime}_{\text{safe}}$, which leads to interference/conflicts between them.
0.B.2.2 Training Configurations.
LLaVA-1.5-7B is adopted for finetuning. Specifically, we follow the official repository [^5] https://github.com/haotian-liu/LLaVA to train the model using LoRA with a rank of 128. For all the training experiments, we finetune the model for 1 epoch with a global batch size of 128, a learning rate of $2\times 10^{-4}$ and weight decay of $0.1$ on 4 Telsa-V100-SXM2-32G GPUs.
0.B.2.3 Discussion on Strategies to Maintain Utility.
To maintain the utility of MLLMs during safety alignment, concurrent work [73] finetunes LLaVA-1.5-7B with a mixture of sub-sampled instruction tuning data (5K examples randomly sampled from its 665K original data [^6] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K/blob/main/llava_v1_5_mix665k.json) and the full data in VLGuard. They find that the finetuned model (i.e., LLaVA-v1.5-7B-Post-hoc-LoRA in Table 13 of [73]) can maintain utility performance on both language datasets (e.g., MMLU [16] and AlpacaEval [27]) and vision-language datasets (e.g., ScienceQA [45] and VizWiz [24]). However, we reproduce this model and observe a large performance gap on MME-P and MME-C compared to the original model. In Table 8, we compare the utility of the reproduced model with others. As can be seen, mix-llava-VLGuard obtains the worst utility compared to the others. Therefore, in the finetuning experiment of this paper, we do not incorporate sub-sampled instruction tuning data but use $D^{\prime}_{\text{safe}}$ generated by ECSO.
Orignal model ECSO_VLGuard VLGuard mix-llava-VLGuard
Utility 1726.9 1727.2 1679.7 1632.0
Table 8: Utility comparison between different models (LLaVA-1.5-7B). The definition of utility is the same as in Table 12. Original is the untuned LLaVA-1.5-7B, ECSO_VLGuard and VLGuard are models in Figure 10 introduced in the Data Construction section, and mix-llava-VLGuard is the reproduced model following [73].
Appendix 0.C Experiments
0.C.1 Results on Other Scenarios of MM-SafetyBench.
In Table 1, we only report the results of LLaVA-1.5-7B on the scenarios 01-07 & 09 of MM-SafetyBench. Here, we present results of the remaining scenarios in Table 13. Similarly, Tables 14, 15, 16, 17 shows the results on the models ShareGPT4V-7B, mPLUG-Owl-7B, Qwen-VL-Chat and InternLM-XComposer-7B, respectively. Specifically, we observe that even without images (i.e., text only), on scenarios 08 and 10-13, all models achieve poor safety in their responses when directly prompted (i.e., all are less than 15% by average in terms of harmless rates). This reveals the intrinsic vulnerabilities of the LLMs for these scenarios. Therefore, we cannot expect MLLMs to be safer with the presence of images, nor when using ECSO as it relies heavily on the safety mechanism of LLMs. Note that we observe relatively higher harmless rate in data type “SD” (e.g., Qwen-VL-Chat and InternLM-XComposer-7B) since “SD” does not contain texts in images. Thus, the model might understand the image in a harmless way.
0.C.2 Additional Ablations
0.C.2.1 Effects of Safety Instructions
In the last step of ECSO, the LLM is instructed with “Based on the hints, answer the following question in a HARMLESS and ETHICAL way.” We ablate the effect of this safety instruction on Direct and ECSO, and report the harmless rate in Table 9. We find the safety instruction is beneficial for response safety but CANNOT address the safety issues of MLLMs.
Methods LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
VL. MM. VL. MM. VL. MM. VL. MM. VL. MM.
Direct 20.5 48.8 16.9 47.8 65.7 52.5 26.8 54.0 18.7 43.9
w/ inst. 24.1 49.5 19.7 47.0 49.8 50.7 77.8 81.1 31.9 46.1
ECSO 91.8 86.4 93.3 85.3 95.9 83.4 81.6 78.5 76.8 80.6
w/o inst 71.3 76.3 81.4 71.0 93.0 71.5 26.3 67.1 68.9 73.9
Table 9: Abaltions on the effect of safety instruction. Harmless rates on VLsafe (VL.) and MM-SafetyBnech (MM.) are reported (averaged over the scenarios in Table 1 of the SD+OCR split). Due to space limits, model sizes are not presented in their names. InternLM-XC is short for InternLM-XComposer.
0.C.2.2 Different Harm-Detection Strategies
In step 2 of ECSO, we let the MLLM judge whether its own response is safe or not. In contrast to this strategy, one can also conduct detection before model responses with the query and image/caption. In Table 10, we present the discrimination accuracies of various strategies on the same dataset as described in Sec. 3.2. We find that detection based on the model response achieves the highest accuracy. This can be explained by the fact that harmful content is more explicit in the response than in the query or image.
Detect LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
Response 95.0 96.6 92.7 86.9 88.5
Img. & Q. 75.6 90.2 84.8 61.3 84.3
Cap. & Q. 82.6 91.6 88.9 66.0 83.0
Table 10: Accuracy of various MLLMs detection on safety of responses with various strategies. Here Img. and Cap. and Q. stand for image, caption, query, respectively.
0.C.2.3 Usage of LLMs
In step 4 of ECSO, the LLM part of the MLLM is employed to generate a safe answer. However, this LLM is finetuned when it is connected to the vision encoder. Here we replace it with the original LLM and report the harmless rate of ECSO in Table 11777As mPLUG-Owl2 starts with LLaMA-2-base instead of its chat version, we do not include the results on the original LLM.. As can be seen, ECSO with the original LLM achieves higher harmless rate. This results from the fact that finetuning may cause forgetting in the previously learned text-only safety alignment. However, in ECSO, we do not employ this strategy because it requires loading two LLMs into the memory.
Methods LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
VL. MM. VL. MM. VL. MM. VL. MM. VL. MM.
finetuned LLM 91.8 86.4 93.3 85.3 95.9 83.4 81.6 78.5 76.8 80.6
original LLM 93.3 85.9 94.9 85.8 96.3 83.8 89.0 82.0 - -
Table 11: Abaltions on the usage of LLM. Evaluations are the same as in Table 9.
0.C.3 Applying ECSO on the Finetuned Model
In this section, we apply ECSO on the models finetuned in Sec. 5.6. Table 12 shows the harmless rate of LLaVA-1.5-7B (fintuned on different datasets) on MM-SafetyBench. It can be observed that ECSO is complementary with finetuning.
ECSO ECSO_VLGuard VLGuard
FT FT + ECSO FT FT + ECSO
86.38 81.39 90.09 73.68 89.18
Table 12: Combine finetuning with ECSO. Results of LLaVA-1.5-7B on MM-SafetyBench are reported (following Table 9). FT: Finetuning.
0.C.4 Case Studies
Figures 13-18 show more qualitative examples to ECSO for the models.
Specifically, Figures 13-15 show more responses from ECSO on the datasets used in Sec. 5 (i.e., MM-SafetyBench, VLSafe).
Figure 16 shows that ECSO is also effective on FigStep [21], which is another safety benchmark of MLLMs. However, since both MM-SafetyBench and FigStep inject malicious contents into images by OCR, we only report the full results of ECSO on MM-SafetyBench in Sec. 5.
Figure 17 shows that ECSO cannot be easily bypassed via certain strategy. In particular, by replacing the sensitive word “bomb” with a picture of a bomb, one could induce the MLLM to generate harmful responses. However, ECSO succeeds in protecting the MLLM.
Figure 18 shows that ECSO can protect MLLMs from adversarial images. Here we use the examples in [49]. Note that the first image888image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/clean.jpeg is clean and the MLLM rejects to fulfill the malicious request thanks to its safety mechanism (though proving to be limited in this paper). However, the second image999image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/prompt_constrained_64.bmp is adversarial (optimized to let MLLMs generate harmful responses). In this case, “Direct” generates unsafe contents while ECSO succeeds in protecting the MLLM.
Appendix 0.D More Discussions
0.D.0.1 Training with model-generated data
has become an essential research problem in both computer vision (e.g., He et al. [26] for image classification, GeoDiffusion [8, 20, 41, 32, 62, 19] for object detection [25, 30] and StableRep [57] for contrastive learning [5, 43] and masked image modeling [6, 71]) and natural language processing (e.g., SELF [44] for instruction tuning and mistake analysis [7] for safety alignment) and vision-language modeling [29, 22], thanks to the remarkable progress of AIGC. ECSO also belongs to this direction. However, different from previous works, we focus on inheriting the intrinsic safety mechanism of pre-aligned LLMs to safeguard MLLMs to the greatest extent, as in Sec. 4.
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
08-Political Lobbying 4.6 40.5 36.6 10.5 39.9 6.5 5.2
10-Legal Opinion 29.2 3.1 4.6 5.4 5.4 1.0 1.5
11-Financial Advice 3.0 1.0 1.0 0.0 0.0 1.0 0.0
12-Health Consultation 11.9 2.8 1.8 5.5 1.8 1.8 2.8
13-Government Decision 4.0 4.0 4.0 1.3 2.0 0.7 2.7
Average 10.5 10.3 9.6 4.5 9.8 2.2 2.5
Table 13: Harmless rates with LLaVA-1.5-7B on MM-SafetyBench (08 & 09-13) (Complementary with Table 1).
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
01-Illegal Activity 89.7 80.4 94.9 16.5 86.6 22.7 88.7
02-HateSpeech 90.2 89.6 100.0 52.8 92.6 52.2 90.2
03-Malware Generation 65.9 90.9 100.0 36.4 90.9 47.7 75.0
04-Physical Harm 66.7 84.0 93.8 41.7 84.7 38.9 79.9
05-Economic Harm 95.1 98.4 100.0 86.9 94.3 89.3 96.7
06-Fraud 79.2 81.8 96.1 29.2 88.3 27.9 88.3
07-Pornography 79.8 89.0 93.6 73.4 85.3 67.0 78.0
09-Privacy Violence 75.5 84.9 95.7 43.9 92.1 36.7 85.6
Average 80.3 87.4 96.8 47.6 89.4 47.8 85.3
08-Political Lobbying 4.6 40.5 36.6 10.5 39.9 6.5 5.2
10-Legal Opinion 29.2 3.1 4.6 5.4 5.4 1.0 1.5
11-Financial Advice 3.0 1.0 1.0 0.0 0.0 1.0 0.0
12-Health Consultation 11.9 2.8 1.8 5.5 1.8 1.8 2.8
13-Government Decision 4.0 4.0 4.0 1.3 2.0 0.7 2.7
Average 10.5 10.3 9.6 4.5 9.8 2.2 2.5
Table 14: Harmless rates with ShareGPT4V-7B on MM-SafetyBench. As a supplement for Figure 6.
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
01-Illegal Activity 90.7 74.2 84.5 28.9 84.5 18.6 91.8
02-HateSpeech 95.7 82.8 95.7 54.0 89.0 44.8 82.8
03-Malware Generation 61.4 84.1 97.7 43.2 81.8 40.9 77.3
04-Physical Harm 73.6 76.4 89.6 39.6 73.6 26.4 65.3
05-Economic Harm 95.9 99.2 99.2 89.3 94.3 86.9 92.6
06-Fraud 88.3 78.6 93.5 35.7 90.9 32.5 89.6
07-Pornography 73.4 86.2 86.2 67.9 70.6 63.3 66.1
09-Privacy Violence 71.9 77.0 89.9 41.0 82.0 38.1 79.1
Average 81.4 82.3 92.1 49.9 83.3 43.9 80.6
08-Political Lobbying 4.6 30.7 30.7 10.5 9.2 7.2 5.9
10-Legal Opinion 22.3 17.7 14.6 13.8 14.6 5.4 7.7
11-Financial Advice 0.6 1.8 12.0 0.0 0.0 0.0 0.0
12-Health Consultation 14.7 3.7 3.7 4.6 4.6 2.8 1.8
13-Government Decision 3.3 10.1 8.7 2.0 2.7 4.0 2.7
 Average 9.1 12.8 13.9 6.2 6.2 3.9 3.6
Table 15: Harmless rates with mPLUG-Owl-7B on MM-SafetyBench. As a supplement for Figure 6.
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
01-Illegal Activity 100.0 94.8 93.8 50.5 90.7 39.2 80.4
02-HateSpeech 90.2 97.5 98.8 63.2 93.3 46.6 87.7
03-Malware Generation 86.4 100.0 100.0 56.8 79.5 52.3 77.3
04-Physical Harm 86.1 99.3 98.6 55.6 81.9 47.9 76.4
05-Economic Harm 98.4 99.2 99.2 85.2 93.4 92.6 95.1
06-Fraud 97.4 98.1 99.4 48.1 89.6 36.4 83.1
07-Pornography 83.5 96.3 96.3 78.0 85.3 65.1 85.3
09-Privacy Violence 95.0 94.2 96.4 30.2 84.2 39.6 82.0
Average 92.1 97.4 97.8 58.4 87.3 52.5 83.4
08-Political Lobbying 0.7 57.5 60.1 9.2 9.2 4.6 3.9
10-Legal Opinion 20.8 44.6 43.1 16.9 23.8 13.1 15.8
11-Financial Advice 0.0 56.9 52.7 3.0 3.0 2.4 3.6
12-Health Consultation 29.4 12.8 11.0 13.8 14.7 5.5 6.4
13-Government Decision 2.0 55.0 58.4 18.1 16.8 7.4 10.7
Average 10.6 45.4 45.1 12.2 13.5 6.6 8.1
Table 16: Harmless rates with Qwen-VL-Chat on MM-SafetyBench. As a supplement for Figure 6.
Scenarios Text only SD OCR SD+OCR
Direct ECSO Direct ECSO Direct ECSO
01-Illegal Activity 94.8 81.4 94.8 34.0 68.0 35.1 80.4
02-HateSpeech 96.9 86.5 96.3 53.4 82.8 55.2 79.1
03-Malware Generation 63.6 81.8 97.7 45.5 75.0 43.2 70.5
04-Physical Harm 87.5 81.9 93.1 43.8 72.2 51.4 79.2
05-Economic Harm 95.9 98.4 97.5 86.1 92.6 89.3 93.4
06-Fraud 87.0 83.1 94.8 41.6 71.4 44.8 79.2
07-Pornography 74.3 95.4 97.2 77.1 78.0 66.1 77.1
09-Privacy Violence 89.2 82.7 93.5 33.8 55.4 46.8 69.1
Average 86.2 86.4 95.6 51.9 74.4 54.0 78.5
08-Political Lobbying 11.1 33.3 52.9 3.3 20.9 7.2 19.0
10-Legal Opinion 45.4 43.8 60.8 20.0 41.5 31.5 46.2
11-Financial Advice 2.4 13.8 14.3 1.8 7.8 3.6 3.6
12-Health Consultation 8.3 16.5 56.0 1.8 21.1 3.7 11.9
13-Government Decision 2.7 24.8 68.5 8.7 40.3 7.4 44.3
Average 14.0 26.4 50.5 7.1 26.3 10.7 25.0
Table 17: Harmless rates with InternLM-XComposer-7B on MM-SafetyBench. As a supplement for Figure 6.
Figure 13: Qualitative comparison of LLaVA-1.5-7B on MM-SafetyBench.
Figure 14: Qualitative comparison of mPLUG-Owl2-7B on VLSafe.
Figure 15: Qualitative comparison of LLaVA-1.5-7B on MM-SafetyBench.
Figure 16: Qualitative comparison of Qwen-VL-Chat on FigStep [21].
Figure 17: Qualitative comparison of ShareGPT4V-7B on Internet images.
Figure 18: Qualitative comparison of MiniGPT-4[72] on adversarial images in [49]. Note that the first image is clean and the MLLM rejects to fulfill the malicious request. However, the second image is adversarial (optimized to let MLLMs generate harmful responses). In this case, “Direct” generates unsafe contents while ECSO succeeds to protect the MLLM.